OpenStudy (anonymous):

The UAV is flying at a speed of 13 meters per second in the direction of the fire. Suppose the altitude of the UAV is now 800 meters. From its position as point F, how many minutes, to the nearest tenth, will it take the UAV to be directly over the fire?

OpenStudy (anonymous):

depends on which UAV it is, MQ-9B Reaper, MH-2 Robot, RQ-4A Global Hawk? ;) I'm kidding. I'm a military guy that loves this stuff. ANYWAYS! Well your UAV is traveling at 13 meters per second, which is 22.3694 MPH, or which in military terms is: 19.43847715 knots. From the position of point F, it would be 1.8. 13/1 = 1600/x. 123s/60min. =2.05min. and x =123s. So if you relate that down your answer would be 1.8. Hope this helps, have a nice day! Also cute girl in the picture ;) Tag me on anymore math questions you need help with!

OpenStudy (anonymous):

Did this help you alright? If not I can attempt to explain it a little more.

OpenStudy (anonymous):

It did thank you!

OpenStudy (anonymous):

@camila1315 good to hear! :) Glad I could help, no problem. Have a nice day!